OpenStudy (anonymous):

Write an equation of a line whose y-intercept is 2 and whose slope is -1. y = -x + 2 y = x -1 y = -x -2 y = x + 1

OpenStudy (anonymous):

B

jimthompson5910 (jim_thompson5910):

Hint: Use the idea that y = mx+b has a slope of m and a y-intercept of b

OpenStudy (foolaroundmath):

a general equation with slope m and y-intercept c is given as y = mx +c Can you try now ?

OpenStudy (anonymous):

y = -x + 2

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

oh okay Im gonna try it